OpenStudy (anonymous):

what is the answer to 48/2 (9+3)

OpenStudy (anonymous):

the answer is 288 48/2(9+3) *parenthesis first 9+3=12 48/2*12 *Pemdas! 48/2=24 24*12= 288

OpenStudy (anonymous):

thank you, i've been arguing that all day, but some people are too stupid to understand regardless of how much proof i show them online